OpenStudy (anonymous):

I need to solve this equation. 13c+24=13+10c

OpenStudy (anonymous):

3c=-11 3c/3=-11/3 c=-11/3

OpenStudy (highschoolmom2010):

to solve this you need to isolate the variable